# references.bib Discovering Dynamic Symbolic Policies with Genetic Programming

### v2

@@ Abstract
Artificial intelligence (AI) techniques are increasingly being applied to solve control problems. However, control systems developed in AI are often black-box methods, in that it is not clear how and why they generate their outputs. A lack of transparency can be problematic for control tasks in particular, because it complicates the identification of biases or errors, which in turn negatively influences the user’s confidence in the system. To improve the interpretability and transparency in control systems, the black-box structure can be replaced with white-box symbolic policies described by mathematical expressions. Genetic programming offers a gradient-free method to optimise the structure of non-differentiable mathematical expressions. In this paper, we show that genetic programming can be used to discover symbolic control systems. This is achieved by learning a symbolic representation of a function that transforms observations into control signals. We consider both systems that implement static control policies without memory and systems that implement dynamic memory-based control policies. In case of the latter, the discovered function becomes the state equation of a differential equation, which allows for evidence integration. Our results show that symbolic policies are discovered that perform comparably with black-box policies on a variety of control tasks. Furthermore, the additional value of the memory capacity in the dynamic policies is demonstrated on experiments where static policies fall short. Overall, we demonstrate that white-box symbolic policies can be optimised with genetic programming, while offering interpretability and transparency that lacks in black-box models.
@@ 1 Introduction
In contrast to black-box models, white-box models allow for easy inspection of their internal mechanisms [adadi2018peeking]. However, white-box models show an accuracy-interpretability trade-off and there are currently no general training methods to produce white-box approaches that can compete with black-box approaches in terms of efficiency and accuracy [loyola2019black, puiutta2020explainable]. Most white-box approaches opt for post-training model distillation to produce a transparent model that learns the input-output mapping of a black-box model [gou2021knowledge]. Disadvantages of model distillation are that it requires a two-phase learning process and information loss could occur in the distillation phase [ras2022explainable]. Therefore, it may be more desirable to learn white-box models from the start. When training an inherently white-box model, interpretability of decisions already occurs at intermediate stages, showing transparent learning, and outputs are directly dependent on explainable features [ras2022explainable].
@@
One way to learn white-box control systems is to discover symbolic policies [landajuela2021discovering, gou2021knowledge]. Symbolic policy learning is a relatively new field, but it is closely related to symbolic regression [bongard2007automated]. Most applications of symbolic regression focus on discovering the equations that govern unknown dynamics [cao2000evolutionary, sakamoto2001inferring, iba2008inference, schmidt2009distilling, quade2016prediction]. In contrast, symbolic policy learning focuses on discovering the control laws that can optimally control the system it interacts with. Symbolic policies are described by symbolic expressions, consisting of mathematical operators and relevant state variables. Similar to symbolic regression [bongard2007automated], symbolic policy learning does not require a predefined model structure, therefore it is less affected by human bias or knowledge gaps compared to other data-driven approaches. Besides being interpretable, symbolic policies have other advantages such as that the policies may capture rules that generalise better to other settings due to the bias-variance trade-off, and learned equations tend to be more efficient than neural networks in terms of parameter count.
Nonetheless, there are several challenges that make it difficult to learn symbolic expressions. The search space of combinations of operators and variables is extremely large and the structure of a mathematical expressions is non-differentiable, meaning that nonlinear function approximators such as neural networks cannot be easily applied. Furthermore, current methods of symbolic policy learning focus on static policies, and can rarely scale to multiple control dimensions. Dealing with partial observability and varying environment settings are however big challenges in control problems, which requires a policy to have memory to accomplish its goals [heess2015memory]. Therefore it is important to learn dynamic symbolic policies to improve the robustness of the policies.
To address these challenges, we introduce a new evolutionary approach to optimise high-performing white-box models to solve control problems. Evolutionary algorithms have achieved many successes in control and reinforcement learning tasks. Work in [salimans2017evolution] and [such2017deep] showed that genetic algorithms and evolution strategies are powerful alternatives for gradient-based approaches for optimisation of weights of neural networks. NEAT [stanley2002evolving] extends genetic algorithms by evolving neural network architectures from scratch while learning the weights simultaneously. In contrast to these methods, we propose to use genetic programming to learn the structure of dynamical systems that directly implement effective control policies. To this end, we embrace genetic programming (GP). This is an evolutionary approach which offers a gradient-free method to learn computer programs [koza1994genetic]. Therefore, GP can also be used to discover symbolic policies for control problems [hein2018interpretable].
@@
A GP approach is developed to learn both static and dynamic symbolic policies to solve control problems. Formally, static policies learn a fixed mapping from observations to control outputs (Figure 1A). However, this approach falls short when the environment is volatile (changes across time) or partially observable (providing limited information about the state of the environment). In this case, we are in need of dynamic control systems that integrate past observations to make optimal decisions (Figure 1B). In the dynamic case, the GP learns a symbolic representation of a state equation, describing the flow of a differential equation that takes observations as input and generates controls as output. Here, the latent state of the underlying differential equations functions as a memory system to capture long term dependencies in the data, which could improve the policy’s robustness and generalisability [heess2015memory]. This is similar to the use of neural differential equations to estimate parameterized dynamical systems. However, in contrast to neural differential equations we learn a symbolic, interpretable, representation of the state equation that implements the policy.
When symbolic policies with higher dimensionality or with a latent space are optimised, the trees have to be evolved simultaneously, which complicates optimisation of the trees due to the large search space. Our method is extended to improve the efficiency of traversing the search space to discover high-performing static and dynamic symbolic policies consisting of multiple trees. Hence, in the case of dynamic policies, GP learns a symbolic expression for every dimension of the latent state and the readout function (Figure 1C).
@@
Using this approach, symbolic policies are discovered for benchmark linear and non-linear environments as well as a simulation of an industrial problem. Dynamic policies are successfully evolved in environments with partial observability and varying parameter settings whereas static policies obtain limited performance in these settings. High-performing symbolic policies are still discovered when the control dimension increases, showing that our method is scalable. Furthermore, we find that the evolved symbolic policies obtain comparable performance as the black-box baseline, given by neural differential equations [chen2018neural]. The key contribution of our work is a method for discovering interpretable high-performing static and dynamic symbolic policies with GP without requiring human expertise.
@@ 2 Methods
The goal of this paper is to find static and dynamic symbolic policies that obtain high performance in control problems. The general setup in our experiments contains an environment coupled with a static policy (Figure 1A) and a dynamic policy (Figure 1B), which is explained in more detail in this section. Afterwards, the specifics of the GP algorithm and how we extended GP to learn multiple trees efficiently are described. In subsequent subsections the baselines and experiments we performed are detailed. The code including genetic programming implementation, environments and baselines can be found at: https://github.com/sdevries0/MultiTreeGP.
@@ 2.1 Environment model
To evaluate a symbolic policy proposed by GP, a coupled system consisting of the policy and an environment is simulated, as described in Figure 1A and Figure 1B. The environment is modeled by a state , following the dynamics of a stochastic differential equation
where is the control input, is the environment’s state equation, is a multivariate Wiener process and determines how the process noise influences the dynamics. The policies are evaluated on a batch of trajectories, with a batch size of 32. In every trajectory, the initial conditions is randomly sampled, and if applicable a target state and environment parameters are sampled.
The controller does not have direct access to , but instead receives observations about the state , determined by function and possibly corrupted by noise. In our experiments, the noisy observations are generated by
where is a readout matrix and is Gaussian observation noise. When relevant, the policy receives a target state that should be achieved in the environment.
@@ 2.2 Control model
The control model is used to map observations to controls via a readout mechanism. We consider controllers that implement static policies, as conventionally used, as well a controllers that implement dynamic policies, where the control depends on a history of observations.
@@
In case of static policies, we represent the policy as a function , which computes the control signal with only the observations and potential target states as input. This setup matches related approaches for symbolic policy discovery [hein2018interpretable, landajuela2021discovering, guo2024efficient] and is given by
@@
In case of dynamic policies, the policy becomes dependent on a dynamically evolving latent state, modeled by a system of ordinary differential equations, according to
@@ 2.4 Genetic programming
The symbolic policies are evolved via genetic programming. This is achieved by representing the functions and as tree-structured computational graphs. GP is a variant of evolutionary algorithms that focuses on learning computer programs [koza1994genetic]. A population of solutions is evolved to satisfy a certain goal or task through stochastic optimisation. In GP, individuals are represented by a parse tree that is constructed from predefined sets of function nodes and leaf nodes. Function nodes cover mathematical operators and functions, while leaf nodes describe variables and constants. Individuals representing solutions that perform better will produce more offspring than individuals that have worse performance. GP is largely defined by two hyperparameters. The first hyperparameter is the number of generations, which determines how many iterations of evolution are performed. The second hyperparameter is the population size, which defines the number of candidates that are evolved at every generation. See Algorithm 1 (Appendix 2.4) for an overview of the GP algorithm. The GP algorithm consists of initialisation, evaluation and reproduction steps, as described in more detail in the following sections.
@@ 2.4.2 Evaluation
To evaluate an individual, its tree is transformed into a program of desired functionality and tested on a problem. The performance of the individual is expressed with a fitness score, computed with a fitness function. The fitness scores are used to select individuals for reproduction. After evaluation, parents are selected for reproduction using tournament selection [fang2010review].
@@ 2.4.3 Reproduction
In the GP, the new population is built with new candidate solutions that are obtained with crossover and mutation. In crossover the genotype of two individuals are combined to produce new trees. Fitter individuals are more likely to be selected for reproduction, therefore the offspring is built from promising tree structures. In both parents a random node from the tree is selected, after which these nodes and their subtrees are swapped. An example of the crossover operator is shown in Figure 1D. Uniform crossover is a variant of the standard crossover where multiple nodes can be swapped between a pair of trees without swapping full subtrees [spears1991study]. Uniform crossover improves the locality of the crossover operation, but both types of crossover can produce fit individuals and are therefore both included in our implementation.
Mutation was originally not integrated into GP [koza1994genetic], but mutation has shown to be a useful addition to the algorithm [luke1997comparison]. Mutation is applied to a single individual to generate one offspring and is applied instead of crossover in our implementation. Many aspects of trees can be mutated, for example changing, deleting or adding operators, replacing subtrees or changing variables or constants. Figure 1E shows an example of mutation, in which an operator is changed to a different type of operator. Besides crossover and mutation, the new population also consists of randomly generated individuals and the fittest individuals from the previous generation.
The number of trees that has to be optimised differs for the static and dynamic policies. The static policies should include a tree for every control input the problem requires. These trees also have to be evolved in the dynamic policies, but additional trees are optimised for every latent state variable. We chose to fix the size of the latent state in the dynamic policies and allow the readout function to learn how many latent states to use. To learn both a symbolic latent state and readout function simultaneously in the dynamic policies, the individuals in the GP algorithm are extended to learn multiple trees [cao2000evolutionary]. During reproduction, crossover and mutation are not applied to every tree in the individual to be able to make small jumps through the search space. Crossover is only applied to trees in the same position in two parents, and a third type of crossover is introduced in which only complete trees are swapped between individuals.
@@ 2.4.4 Regularisation
Regularisation is necessary in GP to produce interpretable trees and improve the efficiency of the optimisation. To regularise the size and complexity of the reproduced trees, a punishment relative to the number of nodes in a tree is added to the fitness [luke2006comparison]. Additionally, the depth of the trees is limited to seven, and during initialization the initial depth is set to four. To introduce more compact trees in the population, a small part of the population is simplified mathematically at every generation [hooper1996improving]. To improve diversity in the population during evolution, multiple subpopulations are evolved independently [fernandez2003empirical]. By configuring the hyperparameters in each subpopulation differently, exploration and exploitation is enforced explicitly [herrera2000gradual]. Finally, dynamic policies with constant readout functions that only consist of targets or scalars and ignore the latent state often form local minima because they offer stable solutions. However such policies will never come close to high-performing policies. To improve the efficiency of the evolutionary search, every readout function of a dynamic policy that is reproduced or sampled will have at least one reference to a latent variable in its readout function to avoid such local minima, or else it is rejected for the next population.
@@ 2.5 Baselines
The symbolic policies generated with GP are compared with two baselines. The baselines are evaluated with the same population size and number of generations for every experiment.
The first baseline is random search of dynamic symbolic policies, which randomly samples a set of trees for each candidate, with candidates having the same model structure as the dynamic policies (Figure 1C). The random search policies have the same number of trees in the latent state and are built from the same set of functions and leaf nodes as used in GP. The maximum depth of trees found with random search is fixed to seven. Random search is included as a baseline to demonstrate if GP can find better solutions with less computation than random sampling, therefore showing the additional value of the genetic operators during optimisation.
@@
with and . In this case, model parameters are optimised using covariance matrix adaption evolution strategies [hansen2016cma]. Evolutionary optimisation of weights in neural networks was shown to be a powerful alternative for gradient-based training in reinforcement learning tasks [salimans2017evolution]. In all experiments, the dimension of was set to five for the NDE baseline. If policies discovered with GP can obtain similar fitness scores as the black-box NDE, the symbolic policies are said to be high-performing.
@@ 2.6 Environments
Our method and the baselines are evaluated on three different environments, as shown in Figure 2. The first environment is the stochastic harmonic oscillator, which is a linear environment that can be solved with a linear controller. The policies are tested in a noisy setting, a partial observable setting where only a subset of the state variables is observed, and a setting where the environmental parameters vary. Here, an analytically optimal control policy, linear quadratic Gaussian (LQG) control, is used as an optimal baseline for comparison. The second environment is the acrobot swing up task with bounded continuous action space. This is a non-linear environment that requires a non-linear controller to perform the swing up. The policies solve the acrobot under noise and partial state observability, and in a third experiment the policies output two control signals. The third environment is the continuous stirred tank reactor (CSTR). The CSTR is a non-linear industrial system, which shows how well GP can find symbolic policies in more realistic settings. The parameters of the CSTR are varied in every trial. The chosen hyperparameter setting of the GP algorithm for each experiment are presented in Table 1 (Appendix 2.4. The choice of the function sets are described in the following subsections.

### v3

@@ Abstract
Artificial intelligence techniques are increasingly being applied to solve control problems, but often rely on black-box methods without transparent output generation. To improve the interpretability and transparency in control systems, models can be defined as white-box symbolic policies described by mathematical expressions. While current approaches to learn symbolic policies focus on static policies that directly map observations to control signals, these may fail in partially observable and volatile environments. We instead consider dynamic symbolic policies with memory, optimised with genetic programming. The resulting policies are robust, and consist of easy to interpret coupled differential equations. Our results show that dynamic symbolic policies compare with black-box policies on a variety of control tasks. Furthermore, the benefit of the memory in dynamic policies is demonstrated on experiments where static policies fall short. Overall, we present a method for evolving high-performing symbolic policies that offer interpretability and transparency, which lacks in black-box models.
@@ 1 Introduction
In contrast to black-box models, white-box models allow for easy inspection of their internal mechanisms [adadi2018peeking]. However, white-box models show an accuracy-interpretability trade-off and there are currently no general training methods to produce white-box approaches that can compete with black-box approaches in terms of efficiency and accuracy [loyola2019black, puiutta2020explainable]. Most white-box approaches opt for post-training model distillation to produce a transparent model that learns the input-output mapping of a black-box model [gou2021knowledge]. Disadvantages of model distillation are that it requires a two-phase learning process and information loss could occur in the distillation phase [ras2022explainable]. Therefore, it may be more desirable to learn white-box models directly. When training an inherently white-box model, interpretability of decisions already occurs at intermediate stages, showing transparent learning, and outputs are directly dependent on explainable features [ras2022explainable].
@@
A promising candidate for white-box control systems are symbolic policies [calegari2020integration, landajuela2021discovering, gou2021knowledge]. Symbolic policies are described by interpretable symbolic expressions, consisting of mathematical operators and state variables. Similar to symbolic regression [bongard2007automated], symbolic policy learning does not require a predefined model structure, therefore it is less affected by human bias or knowledge gaps compared to other data-driven approaches. Additionally, symbolic policies may capture rules that generalise better to other settings, and learned equations tend to be more efficient than neural networks in terms of parameter count [landajuela2021discovering].
Current approaches of symbolic policy learning are limited to learning static policies that have a fixed mapping from observations to control outputs [guo2024efficient, hein2018interpretable, nadizar2024naturally, landajuela2021discovering, gou2021knowledge, shimooka2000generating]. However, extending static policies to real-world scenarios is difficult, because environments can be volatile (change across time) or partially observable (providing noisy and limited information about the state of the environment). Dynamic policies could in principle solve these issues by incorporating memory to integrate past observations [heess2015memory], but it is as of yet unclear how to optimise such policies in closed-loop control tasks, while incorporating interpretability.
Due to the non-differentiable structure of mathematical expressions, gradient-based methods cannot be used for the optimisation of symbolic policies. A viable alternative is genetic programming (GP), an evolutionary algorithm that learn computer programs [koza1994genetic]. GP has seen great successes in symbolic regression problems, which seek to discover interpretable mathematical models describing data [cao2000evolutionary, sakamoto2001inferring, iba2008inference, schmidt2009distilling, quade2016prediction, bongard2007automated, wang2019symbolic]. Therefore we consider GP as a promising approach to optimise dynamic symbolic policies to solve control tasks [hein2018interpretable, nadizar2024naturally].
@@
In the dynamic policies, the GP learns a symbolic representation of a state equation, describing the flow of a differential equation that takes observations as input and generates controls as output. Here, the latent state of the underlying differential equations functions as a memory system to capture long term dependencies in the data, which could improve the policy’s robustness and generalisability [heess2015memory]. This is similar to the use of neural differential equations to estimate parameterized dynamical systems [chen2018neural]. However, in contrast to neural differential equations we learn a symbolic, interpretable representation of the state equation that implements the policy.
@@
In this paper, we introduce a method for evolving dynamic symbolic policies with GP to solve dynamic control tasks (Fig. 1). Using this approach, dynamic symbolic policies are successfully evolved in linear and non-linear environments, under several challenging circumstances such as partial observability and varying environment settings where static policies obtain limited performance. High-performing symbolic policies are still discovered when the control dimension increases, showing that our method is scalable. Furthermore, we find that the evolved symbolic policies obtain comparable performance as the black-box baseline, given by neural differential equations [chen2018neural]. The key contribution of our work is a method for discovering interpretable high-performing dynamic symbolic policies with GP without requiring human expertise.
@@ 2 Methods
The goal of this paper is to evolve dynamic symbolic policies that obtain high performance in control problems. We compare dynamic symbolic policies with static symbolic policies, which are explained in more detail in this section. Afterwards, the specifics of the GP algorithm and how we extended GP to learn multiple trees efficiently are described. In subsequent subsections the baselines and experiments we performed are detailed. The code including our GP implementation, environments and baselines is available at: https://github.com/sdevries0/MultiTreeGP.
@@ 2.1 Environment model
To evaluate a candidate symbolic policy proposed by GP, a coupled system consisting of the policy and an environment is simulated, as described in Fig. 2a and Fig. 2b. The environment is modeled by a state , following the dynamics of a stochastic differential equation
where is the control input, is the environment’s state equation, is a multivariate Wiener process and determines how the process noise influences the dynamics.
The controller receives observations from the state , determined by function and possibly corrupted by noise. In our experiments, the noisy observations are generated by
where is a readout matrix and is Gaussian observation noise. When relevant, the policy receives a target state that should be achieved in the environment. The policies are evaluated on a batch of trajectories, with a batch size of 32. In every trajectory, the initial conditions are randomly sampled, and if applicable target states and environment parameters are sampled.
@@ 2.2 Control model
Symbolic control systems are conventionally defined as static symbolic policies [landajuela2021discovering, hein2018interpretable, guo2024efficient], which directly map observations to a control signal (Fig. 2a). The policy is represented as a function , which computes the control signal with only the observations and potential target states as input. This setup is given by
@@
We extend this approach to dynamic symbolic policies (Fig. 2b), which include memory by using a latent state. The latent state processes incoming observations, and the readout function maps the latent state to a control signal. The dynamic policies contain expressions for the latent dimension and control dimension (Fig. 2c). The latent state is governed by a system of ordinary differential equation, according to
@@ 2.4 Genetic programming
The symbolic policies are evolved via genetic programming (GP). GP is a variant of evolutionary algorithms that focuses on learning computer programs [koza1994genetic]. A population of solutions is evolved to satisfy a certain goal or task through stochastic optimisation. In our setup, individuals are represented by tree-structured computational graphs, constructed from predefined sets of function nodes and leaf nodes. Function nodes cover mathematical operators and functions, while leaf nodes describe variables and constants. GP is largely defined by two hyperparameters. The first hyperparameter is the number of generations, which determines how many iterations of evolution are performed. The second hyperparameter is the population size, which defines the number of candidates that are evolved at every generation. See Algorithm 1 for an overview of the GP algorithm. The GP algorithm consists of initialisation, evaluation and reproduction steps, as described in more detail in the following sections.
@@ 2.4.2 Evaluation
To evaluate an individual, its tree is transformed into a program of desired functionality and tested on a problem. The performance of the individual is expressed with a fitness score, computed with a fitness function. The fitness scores are used to select individuals for reproduction, where fitter individuals produce more offspring. After evaluation, parents are selected for reproduction using tournament selection [fang2010review].
@@ 2.4.3 Reproduction
In GP, the new population is built with new candidate solutions that are generated with crossover and mutation. In crossover the genotype of two individuals are combined to produce new trees. Fitter individuals are more likely to be selected for reproduction, therefore the offspring is built from promising tree structures. In both parents a random node from the tree is selected, after which these nodes and their subtrees are swapped. An example of the crossover operator is shown in Fig. 2d. Uniform crossover is a variant of the standard crossover where multiple nodes can be swapped between a pair of trees without swapping full subtrees [spears1991study]. Uniform crossover improves the locality of the crossover operation, but both types of crossover can produce fit individuals and are therefore both included in our implementation.
Mutation was originally not integrated into GP [koza1994genetic], but mutation has shown to be a useful addition to the algorithm [luke1997comparison]. Mutation is applied to a single individual to generate one offspring and is applied instead of crossover in our implementation. Many aspects of trees can be mutated, for example changing, deleting or adding operators, replacing subtrees or changing variables or constants. Fig. 2e shows an example of mutation, in which an operator is changed to a different type of operator. Besides crossover and mutation, the new population also consists of randomly generated individuals and the fittest individuals from the previous generation.
The number of trees that has to be optimised differs for the static and dynamic policies. The static policies evolve a tree for every control input the problem requires. These trees also have to be evolved in the dynamic policies, but additional trees are optimised for every latent state variable. We chose to fix the size of the latent state in the dynamic policies and allow the readout function to learn which latent states to use. To learn both a symbolic latent state and readout function simultaneously in the dynamic policies, the individuals in the GP algorithm are extended to learn multiple trees [cao2000evolutionary]. During reproduction, crossover and mutation are not applied to every tree in the individual to be able to make small jumps through the search space. Crossover is only applied to trees in the same position in two parents, and a third type of crossover is introduced in which only complete trees are swapped between individuals.
@@ 2.4.4 Regularisation
Regularisation is necessary in GP to produce interpretable trees and improve the efficiency of the optimisation. To regularise the size and complexity of the reproduced trees, a punishment relative to the number of nodes in a tree is added to the fitness [luke2006comparison]. Additionally, the depth of the trees is limited to seven, and during initialisation the depth is set to four. To introduce more compact trees in the population, a small part of the population is simplified mathematically at every generation [hooper1996improving]. To improve diversity in the population during evolution, multiple subpopulations are evolved independently [fernandez2003empirical]. By configuring the hyperparameters in each subpopulation differently, exploration and exploitation is enforced explicitly [herrera2000gradual]. Finally, dynamic policies with constant readout functions that only consist of targets or scalars and ignore the latent state, often form local minima because they offer stable solutions. However such policies will never come close to high-performing policies. To improve the efficiency of the evolutionary search, every readout function of a dynamic policy that is reproduced or sampled will have at least one reference to a latent variable to avoid such local minima, or else it is rejected for the next population.
@@ 2.5 Baselines
The static and dynamic symbolic policies generated with GP are compared with two baselines. The baselines are evaluated with the same population size and number of generations for every experiment.
The first baseline is random search of dynamic symbolic policies, which randomly samples a set of trees for each policy. The policies have the same model structure as the dynamic policies (Fig. 2c). The random search policies have the same number of trees in the latent state and are built from the same set of functions and leaf nodes as used in GP. The maximum depth of trees found with random search is fixed to seven. Random search is included as a baseline to demonstrate if GP can find better solutions with less computation than random sampling, therefore showing the additional value of the genetic operators during optimisation.
@@
with and . In this case, model parameters are optimised using covariance matrix adaption evolution strategies [hansen2016cma]. In all experiments, the dimension of was set to five for the NDE baseline. If policies discovered with GP can obtain similar fitness scores as the black-box NDE, the symbolic policies are said to be high-performing.
@@ 2.6 Environments
Our method and the baselines are evaluated on three different environments, as shown in Fig. 3. The first environment is the stochastic harmonic oscillator (SHO), which is a linear environment that can be solved with a linear controller. The policies are tested in a noisy setting, a partial observable setting where only a subset of the state variables is observed, and a setting where the environmental parameters vary. Here, an analytically optimal control policy, linear quadratic Gaussian (LQG) control, is used as an optimal baseline for comparison. The second environment is the acrobot swing up task with bounded continuous action space. This is a non-linear environment that requires a non-linear controller to perform the swing up. The policies solve the acrobot under noise and partial state observability. The third environment is the continuous stirred tank reactor (CSTR). The CSTR is a non-linear industrial system, which shows how well GP can find symbolic policies in an industrial setting. The parameters of the CSTR are varied in every trial and the state is partially observed. The chosen hyperparameter setting of the GP algorithm for each experiment are presented in Table 2 (Appendix 2.4). The choice of the function sets are described in the following subsections.